# and quantum information science

### v2

@@ 4.3 Growth of complexity under random circuits
Known barriers such as the natural proof barrier discovered by Razborov and Rudich [70, 71, 72] prevent us from proving superpolynomial lower bounds on the circuit complexity of an explicit Boolean function. The minimal number of elementary gates (such as AND, NAND, and XOR) in a Boolean function’s implementation is called its circuit complexity. A similar obstructions might inhibit us from proving superpolynomial lower bounds on the quantum complexity of quantum states, that is, the minimal number of $2$-local quantum gates required for construction of the target state from a similar reference. One can more easily prove lower bounds on the complexities of states prepared with random circuits as the latter quickly assume properties of the Haar measure, counting arguments about which yield lower bounds on the circuit complexity. In 1949 Claude Shannon [73] observed that the number of Boolean functions $f:\{0,1\}^{n}\to\{0,1\}$ outgrows the number of functions implementable with less than $\sim 2^{n}/n$ gates. A similar counting argument applies to Haar-random unitaries [74].
@@ 4.4 Linear growth of exact circuit complexity
Here, we review a proof of the linear growth of the exact circuit complexity [75] for random circuits. More precisely, we show that the exact circuit complexity grows linearly with the number of layers, up to exponentially depth, with unit probability over the choice of circuits. Moreover, the complexity saturates for exponentially deep circuits, again with unit probability. The exact circuit complexity quantifies the minimal number of gates required to prepare a state or effect a unitary exactly. The preparation and action of the unitary cannot include errors. Consequently, some unitaries arbitrarily close to the identity operator have exact circuit complexities of $e^{\Omega(K)}$, [^8] In this section, we use big-Omega notation, $\Omega(X)$, to mean, roughly, “a term that grows at least as quickly as $X$”. where $K$ denotes the system size, or number of degrees of freedom (qubits).
We now sketch a proof of the conjecture [55] that circuit complexity grows linearly for an exponentially long time, as applied to random circuits and the exact circuit complexity. We will compare the degrees of freedom in two sets of unitaries: the unitaries generated by the depth-$D$ circuits and the unitaries generated by smaller, depth-$D^{\prime}$ circuits. The basic idea echoes the fact that, in $\mathbb{R}^{3}$, a plane has no volume. In terms of a well-behaved probability measure on a manifold $M$, every lower-dimensional submanifold has a probability of 00. Denote by $\mathcal{U}_{R}$ the set of all unitaries that can be generated by $R$ gates each. Fortunately, $\mathcal{U}_{R}$, although not a manifold, turns out to be a semialgebraic set defined by constraints $\{f_{i}\geq 0,g_{i}>0,h_{i}=0\}$, where $f_{i},g_{i},h_{i}:{\rm SU}(D)\to\mathbb{R}$ are polynomials. Every semialgebraic set has a well-defined dimension similar to a manifold’s dimensionality.
Using the constant-rank theorem, we can apply the dimension-comparison logic above to random circuits: let $A\subset\mathcal{U}_{R}$ denote a semialgebraic set with $\dim(A)<\dim(\mathcal{U}_{R})$. If we draw elements of $\mathcal{U}_{R}$ from random circuits, then elements of $A$ appear with probability 00.
One can lower-bound $\dim(\mathcal{U}_{R})$ in multiple ways [75]. Reference [76] contains a particularly simple proof of this fact for circuits in a brickwork architecture. The bound is $\sim R/K$, roughly the number of layers. As a consequence, we find the following lower bound for the exact circuit complexity $\mathcal{C}_{0}(U)$ of a random quantum circuit. The minimal number of gates in an exact circuit implementation of $U$ is
@@ 4.5 Linear growth of approximate circuit complexity
The simple argument of Sec. 4.4 is fine-tuned to the exact circuit complexity. Using combinatorial methods, we can obtain similar results about more operational notions of circuit complexity such as the approximate circuit complexity: the minimal number of gates to approximate a state or unitary up to some fixed error. Below, we present such a combinatorial argument, using approximate unitary designs (defined below) [77, 78, 79]. Unitary designs were first used to lower-bound quantum complexity in [80, 81, 82].
Unitary designs are probability distributions on the unitary group SU$(D)$. One cannot distinguish a unitary $t$-design from the Haar measure using only expectation values of degree-$t$ polynomials. More formally, let $\nu$ denote any probability distribution on SU$(D)$. Denote by $\mathbb{E}_{U\sim\nu}f(U)$ the expectation value of a random variable $f:\mathrm{SU}(D)\to\mathbb{R}$, over unitaries drawn according to $\nu$. We call $\nu$ a unitary $t$-design if
@@
The rough argument above applies to every ensemble of states: almost every state from any ensemble $\nu$ has a circuit complexity $\geq\max_{\psi}\log\big{(}\nu\bm{(}B_{\varepsilon}(\lvert\psi\rangle)\bm{)}\big{)}$. Consider the ensemble of states generated by random circuits. We can lower-bound almost every state’s approximate circuit complexity by upper-bounding $\max_{\lvert\psi\rangle}\{\nu_{K}\bm{(}B_{\varepsilon}(\lvert\psi\rangle)\bm{)}\}$. The design property (14) implies lower bounds on the approximate circuit complexity, via higher moments of the Haar measure [82, 80, 81]. By the above counting argument for Haar-random states, a unitary drawn from a design on SU$(2^{K})$ has, with high probability, an approximate circuit complexity lower-bounded by $\Omega(Kt)$.
We can combine this bound with the quick convergence of random circuits to unitary designs. Brown and Viola provided evidence for fast convergence by analyzing the spectra of moment operators via mean-field techniques [83]. Reference [81] then rigorously showed that brickwork random circuits of depth $T\geq Ct^{9.5}\,[2Kt+\log_{2}(1/\varepsilon)]$ [^9] We use the parameter $T$ to denote the number of layers in a brickwork layout. This parameter is similar to the parameter $n$ presented in Sec. 4.1. This change in notation is to avoid confusion as $n$ is almost exclusively used for the system size in the context of quantum information theory. are $\varepsilon$-approximate unitary $t$-designs, for a constant $C>0$ (whose form the authors calculated) and for all $t\leq O(2^{2K/5})$. An approximate design is a probability distribution that relaxes the equality in Eq. (14). In particular, the notion of approximation we get is surprisingly strong and implies relative errors (also called multiplicative errors) to Haar random unitaries for the outcome probabilities of any quantum experiment that queries the $t$ copies of $U$ in parallel. See Refs. [81, 84] for details. This relaxation of the design property is sufficient to imply a circuit lower bound of $\mathcal{C}_{\delta}(U)\geq\Omega(Kt)$ with probability $1-e^{-\Omega(Kt)}$. In particular, the probability of significant short-cuts is small, not only in the system-size, but also in the number of copies $t$. The exponent $10.5$ of $t$ in this bound was improved to $5+o(1)$ in [85] and Ref. [84] reduced the $t$ dependence to an optimal, linear scaling. More precisely, random circuits are $\varepsilon$-approximate unitary $t$-designs in depth
@@
$\mathcal{C}_{\delta}(U)$ denotes the minimal number of $2$-local unitaries necessary to generate any $V$ that is $\delta$-close to $U$ in operator-norm: $||U-V||\leq\delta$. Notice that the probability of significant short-cuts is negligible. In particular, for very deep circuits, this becomes increasingly crucial as the number of random bits that are required to draw a circuit of depth $T$ scales like $\sim KT$.
Combinatorial arguments based on unitary designs are flexible: the idea of using counting arguments via moment bounds was extended to more operational definitions of complexity, such as the minimal number of gates required to distinguish a state from the maximally mixed state [80]. The proof technique relies on unitary designs, as in  [81, 86, 80, 85]. It offers a second advantage over the dimension-comparison technique of [75, 76]: the bound (15) holds also if we draw the gates from a universal gate set—from a finite set $\mathcal{G}\subset{\rm SU}(4)$ whose gates can approximate every $K$-qubit unitary. Moreover, the techniques of Ref. [84] yield similar results for other circuit layouts as well as randomly chosen layouts.
Similar techniques can be used to study the growth of quantum circuit complexity in models of time-dependent dynamics. For example, a discretized model of Brownian motion converges to approximate designs as quickly as random circuits do [87]. Reference [88, 89] provides evidence for a similar convergence by continuous Brownian dynamics. These continuous versions can be directly formulated in the context of holography [90, 91] as we discuss in Sec. 7.
@@ 4.6 Barriers to proving that quantum complexity grows superpolynomially at late times
Each of the methods discussed above exploits a counting argument. The argument concerns a placeholder quantity that roughly measures the ensemble’s randomness. Finding such a quantity for an ensemble is far easier than finding a similarly growing quantity for individual circuits. In the exact-circuit-complexity proof, the placeholder is the “accessible dimension” $\dim\mathcal{U}_{R}$. In the approximate-circuit-complexity proof, any quantity that signifies the convergence to designs will do, such as frame potentials [82] or the expected overlaps $\mathbb{E}_{V_{1},\ldots,V_{d}}|\langle\psi|V_{d}\cdots V_{1}|0^{K}\rangle|^{2}$ for random $2$-qubit gates $V_{1},\ldots,V_{d}$. This strategy dominates all our proofs of lower bounds on circuit complexity beyond linear depth. Indeed, this limitation might be fundamental as circuit complexity is known to be notoriously difficult to bound.
But these comments apply to individual states. What about time evolution under randomly drawn local Hamiltonians? This scenario is closer to the setting considered by Brown and Susskind [55], who conjectured that the quantum complexity grows linearly for an exponentially long time for a typical time-independent local Hamiltonian. Still, the total number of distinct local time independent Hamiltonians is upper bounded by $2^{\mathrm{poly}(K)}$; so clearly the above counting argument fails to establish superpolynomial lower bounds on the approximate quantum circuit complexity.
The problem of proving lower bounds on the quantum complexity for the time-evolution of local Hamiltonians turns out to be related to the separation of computational-complexity classes: Aaronson and Susskind [53] point out that the circuit complexity of exponentially long-time evolution by local Hamiltonians is superpolynomial if and only if $\mathrm{PSPACE}\not\subset\mathrm{BQP}/\mathrm{poly}$. $\mathrm{PSPACE}$ is the class of all computations that can be performed with polynomial memory. $\mathrm{BQP}/\mathrm{poly}$ is the class of all problems that can be solved with (nonuniform) families of polynomially sized quantum circuits. Separations of complexity classes are notoriously difficult to prove, and few such separations are known. Proving a separation of the kind $\mathrm{PSPACE}\not\subset\mathrm{BQP}/\mathrm{poly}$ appears to require a breakthrough in theoretical computer science. Therefore, even showing the existence of a local Hamiltonian for which the circuit complexity does not saturate after a polynomial time requires major technical advances.
@@ 5 Paradigms for complexity I: Nielsen complexity
[…]
@@
    In Nielsen’s framework, evolution along a continuous trajectory decomposes into a sequence of infinitesimal steps. This decomposition is equivalent to Trotterization, a technique for building a discrete circuit to simulate Hamiltonian time evolution [96]. Hence Nielsen’s complexity is related to an important quantum-computational concept.
@@
    One has substantial freedom in equipping Nielsen’s complexity geometry with a distance measure on the unitary manifold. This freedom mimics the choice of a gate set in the definition of exact and approximate circuit complexity. A priori, it is difficult to compute Nielsen’s complexity because the complexity geometry in interesting cases is generally high-dimensional and highly curved. But different choices of metric can belong to the same equivalence class with the same long-distance behavior. Within an equivalence class, we can sometimes find modestly curved members [95, 97, 19], facilitating the calculation of Nielsen complexity.
@@
    Nielsen complexity, as a continuous quantity, naturally relates to continuous features of QFT and cMERA tensor networks [18, 98], as well as to physical time evolution. In fact, the parallel with cMERA motivated the earliest works on complexity in free QFT [99, 100].
@@ 5.1 Definition of Nielsen complexity
The $Y^{I}(t)$ denote velocities or control functions. They are the components of the vector that is tangent at time $t$ to the trajectory generated by $H(t)$ through the group manifold. Different trajectories represent different ways of generating $U$. We have parameterized each trajectory by $t\in[0,1]$, such that $U(t{=}0)=\mathbbm{1}$ and $U(t{=}1)=U$.
@@
[…]
@@
To define a metric on the Hilbert space, recall first that quantum states should be thought of as equivalence classes of elements that differ by a global phase. So we should first ask, what transformations in $\mathrm{SU}(N)$ leave a particular state invariant up to a phase, as such transformations should induce zero cost. The set of such transformations is the stabilizer of the state. Think of the given state as a basis element of the Hilbert space. Then transformations that only affect the remaining $N-1$ basis elements leave it invariant. Including multiplication by a phase, the stabilizer of a state $\lvert\psi\rangle$ is therefore the maximal proper subgroup $\mathrm{SU}(N-1)\times\mathrm{U}(1)$ of $\mathrm{SU}(N)$.

### v3

@@ 4.3 Growth of complexity under random circuits
Known barriers such as the natural proof barrier discovered by Razborov and Rudich razborov2004feasible, Gowersblog, arora2009computational prevent us from proving superpolynomial lower bounds on the circuit complexity of an explicit Boolean function. The minimal number of elementary gates (such as AND, NAND, and XOR) in a Boolean function’s implementation is called its circuit complexity. A similar obstructions might inhibit us from proving superpolynomial lower bounds (in the system size) on the quantum complexity of quantum states. That is, the minimal number of $2$-local quantum gates required for construction of the target state from a similar reference. One can more easily prove lower bounds on the complexities of states prepared with random circuits as the latter quickly assume properties of the Haar measure, counting arguments about which yield lower bounds on the circuit complexity. In 1949 Claude Shannon shannon1949synthesis observed that the number of Boolean functions $f:\{0,1\}^{n}\to\{0,1\}$ outgrows the number of functions implementable with less than $\sim 2^{n}/n$ gates. A similar counting argument applies to Haar-random unitaries preskill1998lecture.
@@ 4.4 Linear growth of exact circuit complexity
Here, we review a proof of the linear growth of the exact circuit complexity haferkamp2021linear for random circuits. More precisely, we show that the exact circuit complexity grows linearly with the number of layers, up to exponentially depth, with unit probability over the choice of circuits. Moreover, the complexity saturates for exponentially deep circuits, again with unit probability. The exact circuit complexity quantifies the minimal number of gates required to prepare a state or effect a unitary exactly. The preparation and action of the unitary cannot include errors. Consequently, some unitaries arbitrarily close to the identity operator have exact circuit complexities of $e^{\Omega(K)}$, [^10] In this section, we use big-Omega notation, $\Omega(X)$, to mean, roughly, “a term that grows at least as quickly as XX”. where $K$ denotes the system size, or number of degrees of freedom (qubits).
We now sketch a proof of the conjecture Brown:2017jil that circuit complexity grows linearly for an exponentially long time, as applied to random circuits and the exact circuit complexity. We will compare the degrees of freedom in two sets of unitaries: the unitaries generated by the depth-$D$ circuits and the unitaries generated by smaller, depth-$D^{\prime}$ circuits. The basic idea echoes the fact that, in $\mathbb{R}^{3}$, a plane has no volume. In terms of a well-behaved probability measure on a manifold $M$, every lower-dimensional submanifold has a probability of 0. Denote by $\mathcal{U}_{R}$ the set of all unitaries that can be generated by $R$ gates each. Fortunately, $\mathcal{U}_{R}$, although not a manifold, turns out to be a semialgebraic set defined by constraints $\{f_{i}\geq 0,g_{i}>0,h_{i}=0\}$, where $f_{i},g_{i},h_{i}:{\rm SU}(D)\to\mathbb{R}$ are polynomials. Every semialgebraic set has a well-defined dimension similar to a manifold’s dimensionality.
We can apply the dimension-comparison logic above to random circuits: let $A\subset\mathcal{U}_{R}$ denote a semialgebraic set with $\dim(A)<\dim(\mathcal{U}_{R})$. If we draw elements of $\mathcal{U}_{R}$ from random circuits, then elements of $A$ appear with probability 0.
One can lower-bound $\dim(\mathcal{U}_{R})$ in multiple ways haferkamp2021linear. Reference li2022short contains a particularly simple proof of this fact for circuits in a brickwork architecture. The bound is $\sim R/K$, roughly the number of layers. As a consequence, we find the following lower bound for the exact circuit complexity $\mathcal{C}_{0}(U)$ of a random quantum circuit. The minimal number of gates in an exact circuit implementation of $U$ is
@@ 4.5 Linear growth of approximate circuit complexity
The simple argument of Sec. 4.4 is fine-tuned to the exact circuit complexity. Using combinatorial methods, we can obtain similar results about more operational notions of circuit complexity such as the approximate circuit complexity: the minimal number of gates to approximate a state or unitary up to some fixed error. Below, we present such a combinatorial argument, using approximate unitary designs (defined below) dankert2005efficient, gross2007evenly, low2010pseudo. Unitary designs were first used to lower-bound quantum complexity in brandao2021models, brandao2016local, Roberts:2016hpo.
Unitary designs are probability distributions on the unitary group SU(D)(D). One cannot distinguish a unitary $t$-design from the Haar measure using only expectation values of degree-$t$ polynomials. More formally, let $\nu$ denote any probability distribution on SU(D)(D). Denote by $\mathbb{E}_{U\sim\nu}f(U)$ the expectation value of a random variable $f:\mathrm{SU}(D)\to\mathbb{R}$, over unitaries drawn according to $\nu$. We call $\nu$ a unitary $t$-design if
@@
[…]
We can combine this bound with the quick convergence of random circuits to unitary designs. Brown and Viola provided evidence for fast convergence by analyzing the spectra of moment operators via mean-field techniques brown2010convergence. Reference brandao2016local then rigorously showed that brickwork random circuits of depth $T\geq Ct^{9.5}\,[2Kt+\log_{2}(1/\varepsilon)]$ [^11] We use the parameter $T$ to denote the number of layers in a brickwork layout. This parameter is similar to the parameter $n$ presented in Sec. 4.1. This change in notation is to avoid confusion as $n$ is almost exclusively used for the system size in the context of quantum information theory. are $\varepsilon$-approximate unitary $t$-designs, for a constant $C>0$ (whose form the authors calculated) and for all $t\leq O(2^{2K/5})$. An approximate design is a probability distribution that relaxes the equality in Eq. (14). In particular, the notion of approximation we get is surprisingly strong and implies relative errors (also called multiplicative errors) to Haar random unitaries for the outcome probabilities of any quantum experiment that queries the $t$ copies of $U$ in parallel. See Refs. brandao2016local, chen2024incompressibility for details. This relaxation of the design property is sufficient to imply a circuit lower bound of $\mathcal{C}_{\delta}(U)\geq\Omega(Kt)$ with probability $1-e^{-\Omega(Kt)}$. In particular, the probability of significant short-cuts is small, not only in the system-size, but also in the number of copies $t$. The exponent $10.5$ of $t$ in this bound was improved to $5+o(1)$ in haferkamp2022random and Ref. chen2024incompressibility reduced the $t$ dependence to an optimal, linear scaling. More precisely, random circuits are $\varepsilon$-approximate unitary $t$-designs in depth
@@
$\mathcal{C}_{\delta}(U)$ denotes the minimal number of $2$-local unitaries necessary to generate any $V$ that is $\delta$-close to $U$ in operator-norm: $U-V||\leq\delta$. Notice that the probability of significant short-cuts is negligible. In particular, for very deep circuits, this becomes increasingly crucial as the number of random bits that are required to draw a circuit of depth $T$ scales like $\sim KT$.
Very recently, the $K$ dependence of the depth in the generation of approximate $t$-designs was improved to a near optimal scaling of $O(\log(K)\,t\,\mathrm{polylog}(t))$ schuster2024random. While this would further improve the $K$ dependence in Eq. (16), the circuit layout in Ref. schuster2024random is not a random brickwork circuit as discussed in this section and different $t$ require vastly different circuit layouts. Improving the design depth of random quantum circuits in a standard brickwork layout to the same near optimal scaling is currently an open problem.
Combinatorial arguments based on unitary designs are flexible: the idea of using counting arguments via moment bounds was extended to more operational definitions of complexity, such as the minimal number of gates required to distinguish a state from the maximally mixed state brandao2021models. The proof technique relies on unitary designs, as in brandao2016local, brandao2016efficient, brandao2021models, haferkamp2022random. It offers a second advantage over the dimension-comparison technique of haferkamp2021linear, li2022short: the bound (15) holds also if we draw the gates from a universal gate set—from a finite set $\mathcal{G}\subset{\rm SU}(4)$ whose gates can approximate every $K$-qubit unitary. Moreover, the techniques of Ref. chen2024incompressibility yield similar results for other circuit layouts as well as randomly chosen layouts.
Similar techniques can be used to study the growth of quantum circuit complexity in models of time-dependent dynamics. For example, a discretized model of Brownian motion converges to approximate designs as quickly as random circuits do onorati2017mixing. Reference jian2023linear, Guo:2024zmr provides evidence for a similar convergence by continuous Brownian dynamics. These continuous versions can be directly formulated in the context of holography Magan:2024aet, Magan:2025hce as we discuss in Sec. 7.
@@ 4.6 Barriers to proving that quantum complexity grows superpolynomially at late times
Each of the methods discussed above exploits a counting argument. The argument concerns a placeholder quantity that roughly measures the ensemble’s randomness. Finding such a quantity for an ensemble is far easier than finding a similarly growing quantity for individual circuits. In the exact-circuit-complexity proof, the placeholder is the “accessible dimension” $\dim\mathcal{U}_{R}$. In the approximate-circuit-complexity proof, any quantity that signifies the convergence to designs will do, such as frame potentials Roberts:2016hpo or the expected overlaps $\mathbb{E}_{V_{1},\ldots,V_{d}}|\langle\psi|V_{d}\cdots V_{1}|0^{K}\rangle|^{2}$ for random $2$-qubit gates $V_{1},\ldots,V_{d}$. This strategy dominates all our proofs of lower bounds on circuit complexity beyond linear depth. Indeed, this limitation might be fundamental as circuit complexity is known to be notoriously difficult to bound.
But these comments apply to individual states. What about time evolution under randomly drawn local Hamiltonians? This scenario is closer to the setting considered by Brown and Susskind Brown:2017jil, who conjectured that the quantum complexity grows linearly for an exponentially long time for a typical time-independent local Hamiltonian. Still, the total number of distinct local time independent Hamiltonians is upper bounded by $2^{\mathrm{poly}(K)}$; so clearly the above counting argument fails to establish superpolynomial lower bounds on the approximate quantum circuit complexity in the system size.
The problem of proving lower bounds on the quantum complexity for the time-evolution of local Hamiltonians turns out to be related to the separation of computational-complexity classes: Aaronson and Susskind susskind2018black point out that the circuit complexity of exponentially long-time evolution by local Hamiltonians is superpolynomial in the system size if and only if $\mathrm{PSPACE}\not\subset\mathrm{BQP}/\mathrm{poly}$. $\mathrm{PSPACE}$ is the class of all computations that can be performed with polynomial memory. $\mathrm{BQP}/\mathrm{poly}$ is the class of all problems that can be solved with (nonuniform) families of polynomially sized quantum circuits. Separations of complexity classes are notoriously difficult to prove, and few such separations are known. Proving a separation of the kind $\mathrm{PSPACE}\not\subset\mathrm{BQP}/\mathrm{poly}$ appears to require a breakthrough in theoretical computer science. Therefore, even showing the existence of a local Hamiltonian for which the circuit complexity does not saturate after a polynomial time requires major technical advances.
@@ 5 Paradigms for complexity I: Nielsen complexity
Nielsen et al. identified the optimal implementation of a unitary transformation $U$, using complexity geometry Nielsen1, Nielsen2, Nielsen3, which we will now explain. Consider the manifold of the unitary group equipped with an inner product on its tangent space that assigns a large norm to “complex” operations and a small norm to “simple” ones. Equivalently, this inner product defines a “complexity metric” on the unitary group manifold. We have to choose what is complex and what is simple – for example, a simple operation may be one that acts on a small number of qubits. Now consider the time evolution operator of a physical system, $U(t)=\mathcal{\overset{\leftarrow}{P}}\exp(-i\int_{0}^{t}dt^{\prime}H(t^{\prime}))$ where $H$ is the Hamiltonian and $\mathcal{\overset{\leftarrow}{P}}$ indicates path ordering. Over the time interval $[0,T]$, the time evolution operator follows a trajectory from the identity $U(0)=\mathbbm{1}$ to $U(T)$. Nielsen et al. defined the complexity of the unitary $U(T)$ as the length of the shortest geodesic from $\mathbbm{1}$ to $U(T)$ on the unitary group manifold equipped with the complexity metric. The length of this geodesic upper bounds the approximate circuit complexity and lower bounds the exact circuit complexity of implementing $U(T)$ as a composition of discrete gates nielsen2010quantum.
@@
    In Nielsen’s framework, evolution along a continuous trajectory decomposes into a sequence of infinitesimal steps. This decomposition is equivalent to Trotterization, a technique for building a discrete circuit to simulate Hamiltonian time evolution Lloyd96. Hence Nielsen’s complexity is related to an important quantum-computational concept.
@@
    One has substantial freedom in equipping Nielsen’s complexity geometry with a distance measure on the unitary manifold. This freedom mimics the choice of a gate set in the definition of exact and approximate circuit complexity. A priori, it is difficult to compute Nielsen’s complexity because the complexity geometry in interesting cases is generally high-dimensional and highly curved. But different choices of metric can belong to the same equivalence class with the same long-distance behavior. Within an equivalence class, we can sometimes find modestly curved members Nielsen3, Brown:2021uov, Brown:2022phc, facilitating the calculation of Nielsen complexity.
@@
    Nielsen complexity, as a continuous quantity, naturally relates to continuous features of QFT and cMERA tensor networks Haegeman:2011uy, Nozaki:2012zj, as well as to physical time evolution. In fact, the parallel with cMERA motivated the earliest works on complexity in free QFT Jefferson:2017sdb, Chapman:2017rqy.
@@ 5.1 Definition of Nielsen complexity
The $Y^{I}(t)$ denote velocities or control functions. They are the components of the vector that is tangent at time $t$ to the trajectory generated by $H(t)$ through the group manifold. Different trajectories represent different ways of generating $U$. We have parameterized each trajectory by $t\in[0,1]$, such that $U(t{=}0)=\mathbbm{1}$ and U(t=1)=UU(t{=}1)=U.
@@
$U(t)=\mathcal{\overset{\leftarrow}{P}}\exp\left(-i\int_{0}^{t}dt^{\prime}H(t^{\prime})\right)\,.$ (18)
@@
To define a metric on the Hilbert space, recall first that quantum states should be thought of as equivalence classes of elements that differ by a global phase. So we should first ask, what transformations in $\mathrm{SU}(N)$ leave a particular state invariant up to a phase, as such transformations should induce zero cost. The set of such transformations is the stabilizer of the state. Think of the given state as a basis element of the Hilbert space. Then transformations that only affect the remaining N1N-1 basis elements leave it invariant. Including multiplication by a phase, the stabilizer of a state $\lvert\psi\rangle$ is therefore the maximal proper subgroup $\mathrm{SU}(N-1)\times\mathrm{U}(1)$ of $\mathrm{SU}(N)$.